can you please show me how to solve 2y to the second=25y+13
can you please show me how to solve 10u to the third-5u to the second=0

Answer by checkley75(3666)

2Y^2=25Y+13
2Y^2-25Y-13=0
(2Y+1)(Y-13)=0
2Y+1=0
2Y=-1
Y=-1/2 ANSWER.
----------------------------------------------------------------
10U^3-5U^2=0 FACTORING OUT U^2 WE GET
U^2(10U-5)=0
U^2=0
U=0 ANSWER
10U-5=0
10U=5
U=5/10
U=1/2 ANSWER.